OpenStudy (anonymous):

When Jim went to The Supreme Steak House, he had a meal costing $32.92. If the tax on the meal was 5%, what was the total cost of the meal? Round to the nearest cent

OpenStudy (anonymous):

tax=5% of 32.92=(5/100)*32.92=$1.646 so, total cost=32.92+1.646=$ 34.566